# Climplicit: Climatic Implicit Embeddings for Global Ecological Tasks

### v1

@@ Abstract
Deep learning on climatic data holds potential for macroecological applications. However, its adoption remains limited among scientists outside the deep learning community due to storage, compute, and technical expertise barriers. To address this, we introduce Climplicit, a spatio-temporal geolocation encoder pretrained to generate implicit climatic representations anywhere on Earth. By bypassing the need to download raw climatic rasters and train feature extractors, our model uses ×$\times 1000$ fewer disk space and significantly reduces computational needs for downstream tasks. We evaluate our Climplicit embeddings on biomes classification, species distribution modeling, and plant trait regression. We find that linear probing our Climplicit embeddings consistently performs better or on par with training a model from scratch on downstream tasks and overall better than alternative geolocation encoding models.
@@ 2.2 Residual Connection for SIREN
This preserves the normal distribution $\mathcal{N}(0,1)$ of $h_{j}$ required for the stability of SIREN (Sitzmann et al., 2020). Figure 1 illustrates ReSIREN and its improved scaling properties compared to SIREN. See Appendix B.2 and F.2 for more details.
@@ 3 Experiments
We assess Climplicit embeddings with linear probing on three macroecological downstream tasks: Biomes classification, species distribution modeling (SDM) and Plant traits regression (see Appendix D for more details). To illustrate the storage- and compute-intensive nature of training that Climplicit aims to mitigate, we train a model from scratch (FS) with a location encoder (Loc) or a CHELSA encoder (CH) for each downstream task. For Loc, we use a 16-layer ReSIREN with the same hyperparameters as Climplicit, but with $[\lambda,\phi]$ as input. For CH, we incorporate CHELSA values from March, June, September, and December, processing them with a 4-layer residual network following Cole et al. (2023). For Loc $+$ CH, we concatenate both embeddings before the probe. Additionally, we compare Climplicit embeddings with those from SatCLIP (Klemmer et al., 2023), Taxabind (Sastry et al., 2025), SINR (Cole et al., 2023), CSP (Mai et al., 2023a), and GeoCLIP (Vivanco Cepeda et al., 2024), which implicitly represent modalities such as satellite imagery, sound, bioclimatic variables, and species occurrence data.
@@ 4 Results
Table 1: Linear probing on diverse geolocation representations. Mean and standard deviation of ten random initializations of the linear probe. Best. Second best.
@@
Climplicit representations surpass all FS ones, except for FS Loc $+$ CH on Biomes, which slightly outperforms Climplicit but with a larger confidence interval, highlighting the high variance of representations learned from scratch on small datasets compared to pretrained representations. These results validate the potential of Climplicit for scenarios with minimal storage, compute, or deep learning knowledge. Furthermore, we observe that other geolocation encoders fall short of Climplicit on all tasks, with the exception of GeoCLIP on SDM, which we attribute to the high resolution of GeoCLIP embeddings. This highlights the relevance of high-resolution climatic representations for the downstream tasks at hand. Further quantitative and qualitative results can be found in Appendix F.1, F.2, and F.3.
@@ 5 Ablation Studies
We analyze the impact of key design choices in Table 2. SIREN removes the residual connections from our ReSIREN blocks, consistently resulting in a performance drop, with $-1.2$ R2 for plant traits regression. Appendix F.2 further highlights the impact of residual connections with increasing model depth. To explore the impact of learning temporal climatic data representations, All- Months is pretrained to reconstruct the concatenation of all monthly CHELSA variables at each location, while March-Only is pretrained exclusively on March data ($m_{i}=3$). Our results support support the choice of making Climplicit a spatio-temporal location encoder rather than encoding only a single month or the entirety of yearly climatic variables. In No H-SIREN we remove the H-SIREN activation function and observe a performance drop on two of our tasks. rec-CHELSA runs linear probing on the reconstructed CHELSA variables instead of the embeddings before the projection layer. As expected, the latent representation better generalizes compared to reconstructed climatic variables. In CHELSA-CLIP, we pretrain our model with a contrastive objective similar to Klemmer et al. (2023) which proves to produce poorer representations compared to directly regressing CHELSA. Whether this approach extends to more complicated regressions or auto-encoding settings for multi-pixel images is left to future work. Finally, ERA5 compares representations learned from monthly ERA5 (Dee et al., 2011) instead of CHELSA. We find that the ×$\times 30$ higher resolution of CHELSA ($30$ versus $900$ arcsec) allows for learning richer representations for our marcoecological tasks.
Table 2: Ablation analysis with linear probing. Mean and standard deviation of ten random initializations of the linear probe. Best.
@@ 6 Discussion and Conclusion
We introduced Climplicit, a spatiotemporal encoder that produces implicit climatic representations anywhere on Earth, aimed at lowering the storage, compute, and machine learning expertise requirements for downstream practitioners. Our model hinges on ReSIREN, a residual extension of SIREN that allows to learn high-resolution representations. Our Climplicit model occupies $99.9\%$ fewer memory than the CHELSA dataset it was trained on, while producing representations that perform on par or better than from-scratch training on several macroecological tasks. Potential limitations of Climplicit embeddings are their implicit nature and lower resolution, compared to the original climatic raster dataset. Finally, in Appendix E, we outline how Climplicit could be extended to provide embeddings for future climatologies under different climate change scenarios, in turn allowing future predictions for the downstream tasks.
@@ Appendix F Extended Results
Table 3: Linear probing on diverse geolocation representations for non-ecological downstream tasks. Mean and standard deviation of ten random initializations of the linear probe. Best. Best pretrained geolocation representation.
@@ F.1 Non-Ecological Downstream Tasks
F.2 Scaling of ReSIREN
@@ F.3 Qualitative Results
Figure 7: Visualization of the species distribution for Quercus ilex (Stone oak) after probing the ReSIREN networks pretrained with different numbers of layers. All distributions are for the month of September. Blue to yellow signifies the predicted probability and green signifies the occurrences of the oak in September. It shows that with increasing depth, the level of detail also increases, though at the cost of increasing extrapolation (e.g. Skåne county in Sweden). We also present the distribution learned by a 32 layer SIREN.

### v2

@@ Abstract
Deep learning on climatic data holds potential for macroecological applications. However, its adoption remains limited among scientists outside the deep learning community due to storage, compute, and technical expertise barriers. To address this, we introduce Climplicit, a spatio-temporal geolocation encoder pretrained to generate implicit climatic representations anywhere on Earth. By bypassing the need to download raw climatic rasters and train feature extractors, our model uses ×$\times 3500$ less disk space and significantly reduces computational needs for downstream tasks. We evaluate our Climplicit embeddings on biomes classification, species distribution modeling, and plant trait regression. We find that single-layer probing our Climplicit embeddings consistently performs better or on par with training a model from scratch on downstream tasks and overall better than alternative geolocation encoding models.
@@ 2.2 Residual Connection for SIREN
This preserves the normal distribution $\mathcal{N}(0,1)$ of $h_{j}$ required for the stability of SIREN (Sitzmann et al., 2020). Figure 1 illustrates ReSIREN and its improved scaling properties compared to SIREN. See Appendix B.2 and F.4 for more details.
@@ 3 Experiments
We assess Climplicit embeddings with single-layer probing on three macroecological downstream tasks: Biomes classification, species distribution modeling (SDM) and Plant traits regression (see Appendix D for more details). To illustrate the storage- and compute-intensive nature of training that Climplicit aims to mitigate, we train a model from scratch (FS) with a location encoder (Loc) or a CHELSA encoder (CH) for each downstream task. For Loc, we use a 16-layer ReSIREN with the same hyperparameters as Climplicit, but with $[\lambda,\phi]$ as input. For CH, we incorporate CHELSA values from March, June, September, and December, processing them with a 4-layer residual network following Cole et al. (2023). For Loc $+$ CH, we concatenate both embeddings before the probe. Additionally, we compare Climplicit embeddings with those from SatCLIP (Klemmer et al., 2023), Taxabind (Sastry et al., 2025), SINR (Cole et al., 2023), CSP (Mai et al., 2023a), and GeoCLIP (Vivanco Cepeda et al., 2024), which implicitly represent modalities such as satellite imagery, sound, bioclimatic variables, and species occurrence data.
@@ 4 Results
Table 1: Single-layer probing on diverse geolocation representations. Mean and standard deviation of ten random initializations of the single-layer probe. Best. Second best.
@@
Climplicit representations surpass all FS ones, except for FS Loc $+$ CH on Biomes, which slightly outperforms Climplicit but with a larger confidence interval, highlighting the high variance of representations learned from scratch on small datasets compared to pretrained representations. These results validate the potential of Climplicit for scenarios with minimal storage, compute, or deep learning knowledge. Furthermore, we observe that other geolocation encoders fall short of Climplicit on all tasks, with the exception of GeoCLIP on SDM, which we attribute to the high resolution of GeoCLIP embeddings. This highlights the relevance of high-resolution climatic representations for the downstream tasks at hand. Note in Appendix F.2 Table 4 that using MLP probing lets Climplicit outperform the other methods on all three tasks. Further quantitative and qualitative results can be found in Appendix F.1, F.4, and F.5.
@@ 5 Ablation Studies
We analyze the impact of key design choices in Table 2. SIREN removes the residual connections from our ReSIREN blocks, consistently resulting in a performance drop, with $-1.2$ R2 for plant traits regression. Appendix F.4 further highlights the impact of residual connections with increasing model depth. To explore the impact of learning temporal climatic data representations, All- Months is pretrained to reconstruct the concatenation of all monthly CHELSA variables at each location, while March-Only is pretrained exclusively on March data ($m_{i}=3$). Our results support support the choice of making Climplicit a spatio-temporal location encoder rather than encoding only a single month or the entirety of yearly climatic variables. In No H-SIREN we remove the H-SIREN activation function and observe a performance drop on two of our tasks. rec-CHELSA runs single-layer probing on the reconstructed CHELSA variables instead of the embeddings before the projection layer. As expected, the latent representation better generalizes compared to reconstructed climatic variables. In CHELSA-CLIP, we pretrain our model with a contrastive objective similar to Klemmer et al. (2023) which proves to produce poorer representations compared to directly regressing CHELSA. Whether this approach extends to more complicated regressions or auto-encoding settings for multi-pixel images is left to future work. Finally, ERA5 compares representations learned from monthly ERA5 (Dee et al., 2011) instead of CHELSA. We find that the ×$\times 30$ higher resolution of CHELSA ($30$ versus $900$ arcsec) allows for learning richer representations for our marcoecological tasks.
Table 2: Ablation analysis with single-layer probing. Mean and standard deviation of ten random initializations of the single-layer probe. Best.
@@ 6 Discussion and Conclusion
We introduced Climplicit, a spatiotemporal encoder that produces implicit climatic representations anywhere on Earth, aimed at lowering the storage, compute, and machine learning expertise requirements for downstream practitioners. Our model hinges on ReSIREN, a residual extension of SIREN that allows to learn high-resolution representations. Our Climplicit model occupies $99.9\%$ less memory than the CHELSA dataset it was trained on, while producing representations that perform on par or better than from-scratch training on several macroecological tasks. Potential limitations of Climplicit embeddings are their implicit nature and lower resolution, compared to the original climatic raster dataset. Finally, in Appendix E, we outline how Climplicit could be extended to provide embeddings for future climatologies under different climate change scenarios, in turn allowing future predictions for the downstream tasks.
@@ Appendix F Extended Results
Table 3: Single-layer probing on diverse geolocation representations for non-ecological downstream tasks. Mean and standard deviation of ten random initializations of the single-layer probe. Best. Best pretrained geolocation representation.
@@ F.1 Non-Ecological Downstream Tasks
F.2 MLP Results
Table 4: MLP probing on diverse geolocation representations. Mean and standard deviation of ten random initializations of the linear probe. Best. Second best.
FS Loc 73.1 $\pm$ 3.2 1.64 $\pm$ 0.60 70.6 $\pm$ 0.9
FS CH 66.6 $\pm$ 1.8 3.16 $\pm$ 0.08 74.4 $\pm$ 1.2
FS Loc $+$ CH 75.7 $\pm$ 3.1 3.17 $\pm$ 0.11 76.4 $\pm$ 0.9
SatCLIP 72.3 $\pm$ 0.7 2.80 $\pm$ 0.07 76.9 $\pm$ 0.3
Taxabind 63.6 $\pm$ 1.1 3.04 $\pm$ 0.07 71.4 $\pm$ 0.3
SINR 64.4 $\pm$ 0.8 2.68 $\pm$ 0.07 72.8 $\pm$ 0.7
CSP 64.2 $\pm$ 1.3 3.06 $\pm$ 0.05 72.0 $\pm$ 0.4
GeoCLIP 64.9 $\pm$ 0.6 3.20 $\pm$ 0.06 71.6 $\pm$ 0.3
Climplicit (Ours) 78.2 $\pm$ 0.7 3.70 $\pm$ 0.04 78.6 $\pm$ 0.3
In Table 4, instead of using a single layer as probe, we also calculate results when training a multi-layer perceptron consisting of three hidden layers with 64 neurons on the various embeddings. This is similar to the downstream setup of Klemmer et al. (2023), although the shape of this MLP has not been fine-tuned. We can see that Climplicit outperforms all comparison methods in this setup.
F.3 CHELSA Reconstruction
Figure 3: Analysis of the reconstruction error of CHELSA after the Climplicit pretraining. Top: Absolute error distribution for each climatic variable. Middle: Absolute error distribution for each month. Bottom: Mean absolute error inside 136 by 320 grid cells across the globe.
We samples 100k locations across earth and calculate the absolute error between the original CHELSA values and the reconstruction by Climplicit for each location and month. Note that we calculate this in the Gaussian-normalized space to keep values comparable between the different variables. We also apply log to make the visualizations less sensitive to extreme values.
In Figure 3 we observe that Climplicit reconstructs temperature and radiation the best, but struggles significantly with wind speeds. There is no discernible variation across months. Spatially, the error is concentrated in mountainous areas. As CHELSA is a down sampling of ERA5 data based on a digital elevation model, it follows that the local variation in CHELSA is strongly correlated with elevation changes. Thus the smooth reconstruction by Climplicit performs well in flat areas where CHELSA is equally smooth, but struggles in areas with significant elevation changes. Future work can tackle this issue by oversampling areas with high elevation variability to focus the models capacity there.
F.4 Scaling of ReSIREN
@@ F.5 Qualitative Results
Figure 8: Visualization of the species distribution for Quercus ilex (Stone oak) after probing the ReSIREN networks pretrained with different numbers of layers. All distributions are for the month of September. Blue to yellow signifies the predicted probability and green signifies the occurrences of the oak in September. It shows that with increasing depth, the level of detail also increases, though at the cost of increasing extrapolation (e.g. Skåne county in Sweden). When using MLP probing, we do not observe such extrapolation anymore. We also present the distribution learned by a 32 layer SIREN.